# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 2.4 Action path stochastic decision forests
An important result of this paper, expressed in the next theorem, is that the preceding construction – encompassing a large part of decision problems and games, including very general stochastic versions of it – is well-defined and yields an sdf. This provides the basis for formulating a large class of stochastic decision problems in extensive form – already in this paper it is shown how exogenous information and choices adapted to it, and in the following papers how decision problems can be modelled on that basis.
@@
Under assumptions AP.W0, AP.W1, AP.W2, and AP.W3, the triple $(F,\pi,{\bm{\mathsf{X}}})$ from the current subsection defines a stochastic decision forest over $(\Omega,\mathscr{A})$.
As we will discuss in the follow-up article, for well-ordered $\mathbb{T}$, well-posed decision problems can be defined on this action path sdf, while for the general case, this sdf can turn out to be ill-posed. For this latter case, the action path sdf represents an important first step though, in that a modification of it, the action-reaction sdf, a generalisation of a concept in [1], will restore well-posedness, as will be discussed in the third paper of the present series.
Definition 2.20.
An action path sdf is a stochastic decision forest $(F,\pi,{\bm{\mathsf{X}}})$ over an exogenous scenario space $(\Omega,\mathscr{A})$, that is constructed as above out of a set $W$ satisfying assumptions AP.W0, AP.W1, AP.W2, and AP.W3. $(F,\pi,{\bm{\mathsf{X}}})$ is said to be induced by $W$. $\mathbb{A}$ is called its action space and $\mathbb{T}$ its time axis. If $W$ satisfies assumption AP.W4, then the elements of $I$ are called agents and, for any agent $i\in I$, $\mathbb{A}^{i}$ is called the action space for agent $i$.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be an action path sdf and $W=\bigcup F$. Then $\mathfrak{t}$ is strictly decreasing, that is, for all $x,y\in X$ with $x\supsetneq y$ we have $\mathfrak{t}(x)<\mathfrak{t}(y)$. Moreover, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and all $\omega,\omega^{\prime}\in D_{{\bm{\mathsf{x}}}}$, $\mathfrak{t}({\bm{\mathsf{x}}}(\omega))=\mathfrak{t}({\bm{\mathsf{x}}}(\omega^{\prime}))$.
@@
We conclude this section with examples of action path sdf. This includes the illustrative examples from subsection 2.3 and several typical classes of well-known decision problems. It remains an open question at this point whether any stochastic decision forest can be represented as an action path sdf, that is, so to speak, whether for any sdf all branches of all trees can be synchronised.
@@
The two simple stochastic decision forests from subsection 2.3 can be represented as an action path forest with time $\mathbb{T}=\{0,1\}$.
The proof of this lemma in the appendix shows that the representation of the variant $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ as an action path sdf has to specify a “dummy” action at the terminal node $\{w^{\prime}_{12}\}$, without letting an alternative, that is, an action meaning inaction. This is an artefact of the modelling decision to explicitly include a temporal dimension, as reflected in the action path formulation. [^11] Assumption AP.W1 ensures that moves have a unique time associated to them, but terminal nodes need not. In many cases they do not, as made apparent by the lemmata 2.16 and 2.17 and, most strikingly, their corollary 2.18. But there may be an instant of time that certain branches still “move” at, and others do not and actually identify that instant with later points in time, which appears a bit artificial. Stochastic decision forests are based on first principles and do no include such a dimension. Therefore, apart from being more general and flexible, their decision-theoretic interpretation is much clearer. At the same time, they include action path sdfs which model a structure present in many applications and which can therefore simplify the formal representation, at the cost of possibly introducing artificial phenomena like inactive activity. It may be noted that, in action path sdf, time plays a role complementary to that of random moves: it defines similarity of moves across branches, while random moves define similarity across trees. The random moves in action path sdfs are defined such as to be perfectly compatible with time.
@@
  • 1.
    For $\mathbb{T}=\mathbb{R}_{+}$, singleton $\Omega$, $W=\Omega\times\mathbb{A}^{\mathbb{T}}$, the action path stochastic decision forest yields the game tree of what is called “differential game” in [3, paragraph 2.2.5]. Hence, the action path sdf is a generalisation of this example in several directions: first and foremost, by adding an exogenous scenario space, second, by generalising the time axis, and third, by allowing for much higher flexibility regarding the set of outcomes.
  • 2.
    For $\mathbb{T}=\{0,\dots,N\}$, for $N\in{\mathbb{N}^{\ast}}$, or $\mathbb{T}=\mathbb{N}$ the corresponding action path sdf can serve as a basis describe the classical finite or infinite horizon discrete time, stochastic decision problem (see, e.g., [10]). The possible actions can be specified through the choice of $W$. The case $W=\Omega\times\mathbb{A}^{\mathbb{T}}$ corresponds to the case where at each move, the set of possible actions is $\mathbb{A}$. It is shown in the appendix that this $W$ induces an action path sdf, even for general $\mathbb{T}$.
    Note that in contrast to the traditional “nature” model of stochastic games and decision problems in discrete time, exogenous information cannot be deduced from the order-theoretic properties of the (random) moves. The “nature” player or agent is replaced with a forest of decision trees and a structure of random moves to that we will later attach exogenous information in the form of $\sigma$-algebras.
  • 3.
    $\mathbb{T}$ can also equal more general well-orders: For $\mathbb{T}=\{1-2^{-n}\mid n\in\mathbb{N}\}\cup\{1\}$, the long cheap talk game tree (see [7]) – and general stochastic variants thereof in the sense of sdfs – can be obtained. This example, in case of singleton $\Omega$, is treated in [3].
  • 4.
    Suppose that assumption AP.W4 is satisfied with $\mathbb{A}^{i}=\{0,1\}$ for all $i\in I$. Let $W$ be the set of pairs $(\omega,f)$ where $f\colon\mathbb{T}\to\mathbb{A}$ is component-wise decreasing. It is shown in the appendix that $W$ induces an action path sdf. The corresponding action path sdf is a natural (and at least approximate) candidate for describing stochastic decision problems of timing (alias stopping) (see, e.g., the monographs [40, 33, 19] for the mathematical theory, and [35, 29] for the link to applications in economics and finance).
  • 5.
    Let us consider an example of a forest that becomes shallower towards its crown in some areas of the exogenous scenario space. We use an example from finance, namely the exercise of an American up-and-out option ([27, chapter 26]). Let $P=(P_{t})_{t\in\mathbb{R}_{+}}$ be a continuous stochastic process on $(\Omega,\mathscr{A})$ with strictly positive real values describing the price of a financial security. The initial price is $P_{0}=1$. As long as the price has not reached $2$, the holder of the option can exercise it, but once the price reaches $2$, the option expires irreversibly. This problem can be modelled using the action path stochastic decision forest associated to the following set of outcomes $W$.
    Take $W$ to be the set of all $(\omega,f)$ where $\omega\in\Omega$ and $f\colon\mathbb{R}_{+}\to\{0,1\}$ is decreasing such that, if $f$ takes the value 00, then $t^{\ast}_{f}=\inf\{t\in\mathbb{R}_{+}\mid f(t)=0\}$ satisfies $\max_{t\in[0,t^{\ast}_{f}]}P_{t}(\omega)<2$. It is shown in the appendix that $W$ induces an action path sdf, and that
@@ 3.1 Filtrations in probability theory
Recall that for a measurable space $(\Omega,\mathscr{A})$, the usual interpretation of $\mathscr{A}$ sees its elements as being those subsets of $\Omega$ that are measurable for a given observer or agent. In the context of what we called exogenous scenario spaces, this had essentially the meaning that for all admissible beliefs (alias probability measures) on $\Omega$, probabilities can be computed for these subsets (see subsection 2.1). However, there is a second meaning to “measurable”, ubiquitous in probability theory and many of its applications.
@@
In probability theory, a filtration is a monotone map $\mathscr{F}$ from a non-empty totally ordered set $\mathbb{T}$ modelling time, typically a suborder of the two-sided compactification of $\mathbb{R}$, to the set of sub-$\sigma$-algebras of $\mathscr{A}$, thus mapping any $t\in\mathbb{T}$ to a sub-$\sigma$-algebra $\mathscr{F}_{t}\subseteq\mathscr{A}$ such that for all $t,u\in\mathbb{T}$ with $t\leq u$, $\mathscr{F}_{t}\subseteq\mathscr{F}_{u}$ (see, e.g., [28, chapter 9] or [14, chapter 3]). The customary interpretation refers to an agent being equipped with that filtration $\mathscr{F}$ such that, for all times $t\in\mathbb{T}$, $\mathscr{F}_{t}$ describes the information the agent has at time $t$, or put differently, $\mathscr{F}_{t}$ is the set of events measurable for that agent at time $t$. Of course, this agent is able to compute probabilities for all $A\in\mathscr{A}$, but the elements of $\mathscr{F}_{t}$ are measurable in an even stronger sense.
@@
Namely, one tacitly supposes that there is one realised scenario $\omega\in\Omega$, drawn “at random”, so to speak, but information about it is only revealed to the agent progressively via $\mathscr{F}$, in the following way: For all instants of time $t\in\mathbb{T}$ and each $E\in\mathscr{F}_{t}$ the agent can discern at time $t$ whether it contains the realised scenario or not. In particular, the agent when equipped with a belief in form of a probability measure on $\mathscr{A}$ can derive updated probabilities and expectations by evaluating conditional probabilities and expectations on $E$. Yet, this notion of measurement must be read with some caution. Actually, this measurement capacity of the agent includes logical operations given by the $\sigma$-algebra property, in particular countable logical disjunctions alias unions. Uncountable operations might be excluded however. Moreover, under the belief of the agent, $E$ may have probability zero and hence the evaluation of conditional probabilities and expectations on $E$ may be meaningless ([28, 14]). This can be illustrated by the following examples.
@@
In the discrete setting, where $\mathscr{A}$ is generated by a countable partition of $\Omega$, the interpretation is evident. Then all $\mathscr{F}_{t}$, $t\in\mathbb{T}$, are also generated by countable partitions, and the partition of later instants of time refine the preceding partitions. The members of the partition generating $\mathscr{F}_{t}$ can be thought of as the agent’s information sets regarding the “nature” agent’s past choices at time $t$ [^12] This will be discussed in detail in the follow-up paper. In that discrete case, thus, filtrations are formally and conceptually a special case of order-theoretic rooted forests over the set $\Omega$, as discussed in section 1.
@@
But in general, $\mathscr{F}_{t}$, for some $t\in\mathbb{T}$, may, for instance, contain all the singletons subsets of $\Omega$, without containing all elements of $\mathscr{A}$, let alone all subsets of $\Omega$ – think of $\Omega$ being the unit interval with Borel $\sigma$-algebra $\mathscr{A}$ and $\mathscr{F}_{t}$ being the set of subsets $A\subseteq[0,1]$ that are countable or have countable complement. So the agent would actually be able to measure the realised scenario without discerning some $A\in\mathscr{A}$, say the interval $[0,\frac{1}{2}]$, although it is a union of measurable singletons. The point is that this union is uncountable, and we do not assume agents to be able to perform uncountable logical operations. Hence, the ability to discern the realised scenario does not necessarily imply the ability to discern all events, that is elements of $\mathscr{A}$. Moreover, all but countably many (and in the case of the Lebesgue measure on $[0,1]$ even all) singletons must have probability zero, and so the ability to discern the realised scenario may not be relevant from a decision-theoretic perspective: If the conditional expectation of some interesting quantity, e.g., a payoff, given $\mathscr{F}_{t}$ is computed, it is completely inconclusive to evaluate it on some $E\in\mathscr{F}_{t}$ whose probability is zero, because the conditional expectation is defined almost surely.
@@ 3.2 Exogenous information structures
Based on the preceding argument about exogenous scenario spaces and filtrations, and bearing in mind the way stochastic decision forests are built over exogenous scenario spaces, the following definition is proposed, which is interpreted and analysed in the sequel.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest over an exogenous scenario space $(\Omega,\mathscr{A})$. An exogenous information structure is a family $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ such that for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$:
  1. 1.
    $\mathscr{F}_{\bm{\mathsf{x}}}$ is a sigma-algebra over $D_{\bm{\mathsf{x}}}$ with $\mathscr{F}_{\bm{\mathsf{x}}}\subseteq\mathscr{A}$;
  2. 2.
    for all ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ with ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}^{\prime}$ and every $E\in\mathscr{F}_{\bm{\mathsf{x}}}$, we have $E\cap D_{{\bm{\mathsf{x}}}^{\prime}}\in\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}}$.
@@
Concerning the sense of this definition, we think of any agent as being equipped with some exogenous information structure. For a random move ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, $\mathscr{F}_{\bm{\mathsf{x}}}$ is interpreted as the set of events $E\in\mathscr{A}$ relevant at ${\bm{\mathsf{x}}}$, meaning that $E\subseteq D_{\bm{\mathsf{x}}}$, and that the agent can measure at ${\bm{\mathsf{x}}}$, meaning that the agent can discern whether $E$ contains the realised scenario or not. This second property is to be read in the sense discussed in the previous subsection 3.1. The more straightforward condition (2) ensures that the agent does not forget: if the agent is at ${\bm{\mathsf{x}}}^{\prime}$ and has been at ${\bm{\mathsf{x}}}$ already, and was able to measure $E$ at ${\bm{\mathsf{x}}}$, then this agent can measure $E$ given the domain of ${\bm{\mathsf{x}}}^{\prime}$ also at ${\bm{\mathsf{x}}}^{\prime}$. Note that, if for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, we have $D_{\bm{\mathsf{x}}}=\Omega$, then an exogenous information structure is nothing but a monotone decreasing map $\mathscr{F}\colon{\bm{\mathsf{x}}}\mapsto\mathscr{F}_{\bm{\mathsf{x}}}$ from ${\bm{\mathsf{X}}}$ into the set of sub-sigma-algebras of $\mathscr{A}$, “monotone decreasing” meaning that for all ${\bm{\mathsf{x}}},{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ with ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}^{\prime}$, we have $\mathscr{F}_{\bm{\mathsf{x}}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}}$.
@@
It should be noted that what we have at hand here is an adaptation of the notion of filtrations from probability theory to decision trees. In both cases, the index set is partially ordered. However, in the former case the order is total, whereas in the latter quite the opposite case is true, since it is a decision tree. Moreover, the underlying forest can “thin out” towards the crowns non-uniformly across trees alias exogenous scenarios, so that, in general, domains of random moves must be taken into account.

### v2

@@ 2.4 Action path stochastic decision forests
Definition 2.20.
[…]
If $(F,\pi,{\bm{\mathsf{X}}})$ defines an sdf on $(\Omega,\mathscr{E})$, the word “candidate” is dropped. $\mathbb{A}=\prod_{i\in I}\mathbb{A}^{i}$ is called action space, $\mathbb{T}$ time axis, $I$ the action index set and $\mathbb{A}^{i}$ the $i$-th action space factor of the action path sdf data, and a fortiori, of the induced action path sdf. [^11] Note that $\mathbb{A}$, $\mathbb{T}$, $I$ are part of the data only and, in contrast to $W$, not uniquely determined by the sdf, although “minimal” representatives might be inferred.
Example 2.21.
For $\mathbb{T}=\mathbb{R}_{+}$, singleton $\Omega$, $W=\Omega\times\mathbb{A}^{\mathbb{T}}$, the induced action path sdf candidate $(F,\pi,{\bm{\mathsf{X}}})$, more precisely, $F$ yields the decision tree of what is called “differential game” in [3, Subsection 2.2.5]. Hence, the action path sdf is a generalisation of this example in several directions: first and foremost, by adding an exogenous scenario space, second, by generalising the time axis, and third, by allowing for much higher flexibility regarding the set of outcomes.
Example 2.22.
For $\mathbb{T}=\{0,\dots,N\}$, for $N\in{\mathbb{N}^{\ast}}$, or $\mathbb{T}=\mathbb{N}$ any corresponding action path sdf candidate can serve as a basis for describing the classical finite or infinite horizon discrete-time stochastic decision problem (see, e.g. [10]). The admissible actions can be specified through the choice of $W$. The case $W=\Omega\times\mathbb{A}^{\mathbb{T}}$ corresponds to the case where at each move, the set of possible joint action is given by $\mathbb{A}$. It is discussed in Example 2.26 that these action path sdf data induce an action path sdf, even for general $\mathbb{T}$.
Note that in contrast to the traditional “nature” model of stochastic games and decision problems in discrete time, exogenous information cannot be deduced from the order-theoretic properties of the (random) moves. The “nature” player or agent is replaced with a forest of decision trees and a structure of random moves to that one may attach exogenous information in the form of $\sigma$-algebras as explained in the next section.
$\mathbb{T}$ can also equal more general well-orders: For $\mathbb{T}=\mathbb{N}\cup\{+\infty\}$ with canonical order, that is, essentially, the second smallest infinite ordinal, the long cheap talk game tree (see [7]) – and general stochastic variants thereof in the sense of sdfs – can be obtained. This example, in case of singleton $\Omega$, is treated in [3].
An important result of this paper, expressed in the next theorem, is that the preceding construction – encompassing a large set of decision problems and games in their elementary decision-theoretic structure, including very general stochastic versions of them – is well-defined and yields an sdf. This provides the basis for formulating a large class of stochastic decision problems in extensive form – already in this paper it is shown how exogenous information and choices adapted to it, and in the following papers how decision problems can be modelled on that basis.
@@
Let $(I,\mathbb{A},\mathbb{T},W)$ be action path sdf data on an exogenous scenario space $(\Omega,\mathscr{E})$. Then, the induced action path sdf candidate is an order consistent and maximal stochastic decision forest on $(\Omega,\mathscr{E})$. It surely non-trivial iff for all $\omega\in\Omega$, there are $f,f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ with $f\neq f^{\prime}$ and $(\omega,f),(\omega,f^{\prime})\in W$.
As will be discussed in the follow-up article, for well-ordered $\mathbb{T}$, well-posed decision problems can be defined on this action path sdf, while for the general case, this sdf can turn out to be ill-posed. For this latter case, the action path sdf represents an important first step though, in that a modification of it, the action-reaction sdf, a generalisation of a concept in [1], will restore well-posedness, as will be discussed in the third paper of the present series.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be the action path sdf induced by action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$. Then $\mathfrak{t}$ is strictly decreasing, that is, for all $x,y\in X$ with $x\supsetneq y$ we have $\mathfrak{t}(x)<\mathfrak{t}(y)$. Moreover, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and all $\omega,\omega^{\prime}\in D_{{\bm{\mathsf{x}}}}$, $\mathfrak{t}({\bm{\mathsf{x}}}(\omega))=\mathfrak{t}({\bm{\mathsf{x}}}(\omega^{\prime}))$.
@@
We conclude this section with examples of action path sdf. This includes the illustrative examples from Subsection 2.3 and several typical classes of well-known decision problems. As announced earlier, it remains an open question at this point whether any “relevant” order consistent and maximal stochastic decision forest can be represented as an action path sdf, that is, so to speak, whether for any such sdf all branches of all trees can be synchronised.
@@
The two simple stochastic decision forests from Subsection 2.3 can be represented as an action path sdf with time $\mathbb{T}=\{0,1\}$.
The proof of this lemma in the appendix shows that the representation of the variant $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ as an action path sdf has to specify a “dummy” action at the terminal node $\{w^{\prime}_{12}\}$, without letting an alternative, that is, an action meaning inaction. This is an artefact of the modelling decision to explicitly include a temporal dimension, as reflected in the action path formulation. [^12] Assumption AP.SDF1 ensures that moves have a unique time associated to them, but terminal nodes need not. In many cases they do not, as made apparent by the Lemmata 2.17 and 2.18 and, most strikingly, their Corollary 2.19. But there may be an instant of time that certain branches still “move” at, and others do not and actually identify that instant with later points in time, which appears a bit artificial. Stochastic decision forests are based on first principles and do no include such a dimension. Therefore, apart from being more general and flexible, their decision-theoretic interpretation is much clearer. At the same time, they include action path sdfs which model a structure present in many applications and which can therefore simplify the formal representation, at the cost of possibly introducing artificial phenomena like inactive activity. It may be noted that, in action path sdf, time plays a role complementary to that of random moves: it defines similarity of moves across branches, while random moves define similarity across trees. The random moves in action path sdfs are defined such as to be perfectly compatible with time.
@@
  • Without further assumptions on the data $(I,\mathbb{A},\mathbb{T},W)$, it holds true that if $W=\Omega\times\mathbb{A}^{\mathbb{T}}$, then they are action path sdf data.
  • Suppose that $\mathbb{A}^{i}=\{0,1\}$ for all $i\in I$. Let $W$ be the set of pairs $(\omega,f)$ where $f\colon\mathbb{T}\to\mathbb{A}$ is componentwise decreasing. Then, the quadruple $(I,\mathbb{A},\mathbb{T},W)$ defines action path sdf data. The corresponding action path sdf is a natural (and at least approximate) candidate for describing stochastic decision problems of timing (alias stopping) (see, e.g. the monographs [47, 37, 19] for the mathematical theory, and [41, 33] for the link to applications in economics and finance).
  • Let us consider an example of a forest that becomes shallower towards its crown in some areas of the exogenous scenario space. We use an example from finance, namely the exercise of an American up-and-out option ([31, Chapter 26]). Let $P=(P_{t})_{t\in\mathbb{R}_{+}}$ be a continuous stochastic process on $(\Omega,\mathscr{E})$ with strictly positive real values describing the price of a financial security. The initial price is $P_{0}=1$. As long as the price has not reached $2$, the holder of the option can exercise it, but once the price reaches $2$, the option expires irreversibly. This problem can be modelled using the action path stochastic decision forest associated to the following data.
    Let $I$ be a singleton, $\mathbb{T}=\mathbb{R}_{+}$, and $\mathbb{A}=\{0,1\}$. Take $W$ to be the set of all $(\omega,f)$ where $\omega\in\Omega$ and $f\colon\mathbb{R}_{+}\to\{0,1\}$ is decreasing such that, if $f$ takes the value 00, then $t^{\ast}_{f}=\inf\{t\in\mathbb{R}_{+}\mid f(t)=0\}$ satisfies $\max_{t\in[0,t^{\ast}_{f}]}P_{t}(\omega)<2$. Then, $W$ induces an action path sdf, and
@@
The proofs of these claims can be found in the corresponding part of the appendix.
@@ 3.1 Filtrations in probability theory
Recall that for a measurable space $(\Omega,\mathscr{E})$, the usual interpretation of $\mathscr{E}$ sees its elements as being those subsets of $\Omega$ that are measurable for a given observer or agent. In the context of what we called exogenous scenario spaces, this had essentially the meaning that for all admissible beliefs (alias probability measures) on $\Omega$, probabilities can be computed for these subsets (see Subsection 2.1). However, there is a second meaning to “measurable”, ubiquitous in probability theory and many of its applications.
@@
In probability theory, a filtration is a monotone map $\mathscr{F}$ from a non-empty totally ordered set $\mathbb{T}$ modelling time, typically a suborder of the two-sided compactification of $\mathbb{R}$, to the set of sub-$\sigma$-algebras of $\mathscr{E}$, thus mapping any $t\in\mathbb{T}$ to a sub-$\sigma$-algebra $\mathscr{F}_{t}\subseteq\mathscr{E}$ such that for all $t,u\in\mathbb{T}$ with $t\leq u$, $\mathscr{F}_{t}\subseteq\mathscr{F}_{u}$ (see, e.g. [32, Chapter 9] or [14, Chapter 3]). The customary interpretation refers to an agent being equipped with that filtration $\mathscr{F}$ such that, for all times $t\in\mathbb{T}$, $\mathscr{F}_{t}$ describes the information the agent has at time $t$, or put differently, $\mathscr{F}_{t}$ is the set of events measurable for that agent at time $t$. Of course, this agent is able to compute probabilities for all $E\in\mathscr{E}$, but the elements of $\mathscr{F}_{t}$ are measurable in an even stronger sense.
@@
Namely, one tacitly supposes that there is one realised scenario $\omega\in\Omega$, drawn “at random”, so to speak, but information about it is only revealed to the agent progressively via $\mathscr{F}$, in the following way: For all instants of time $t\in\mathbb{T}$ and each $E\in\mathscr{F}_{t}$ the agent can discern at time $t$ whether it contains the realised scenario or not. In particular, the agent when equipped with a belief in form of a probability measure on $\mathscr{E}$ can derive updated probabilities and expectations by evaluating conditional probabilities and expectations on $E$. Yet, this notion of measurement must be read with some caution. Actually, this measurement capacity of the agent includes logical operations given by the $\sigma$-algebra property, in particular countable logical disjunctions alias unions. Uncountable operations might be excluded however. Moreover, under the belief of the agent, $E$ may have probability zero and hence the evaluation of conditional probabilities and expectations on $E$ may be meaningless ([32, 14]). This can be illustrated by the following examples.
@@
In the discrete setting, where $\mathscr{E}$ is generated by a countable partition of $\Omega$, the interpretation is evident. Then all $\mathscr{F}_{t}$, $t\in\mathbb{T}$, are also generated by countable partitions, and the partition of later instants of time refine the preceding partitions. The members of the partition generating $\mathscr{F}_{t}$ can be thought of as the agent’s information sets regarding the “nature” agent’s past choices at time $t$ [^13] This will be discussed in detail in the follow-up paper. In that discrete case, thus, filtrations are formally and conceptually a special case of order-theoretic rooted forests on the set $\Omega$, as discussed in Subsection 1.
@@
But in general, $\mathscr{F}_{t}$, for some $t\in\mathbb{T}$, may, for instance, contain all the singletons subsets of $\Omega$, without containing all elements of $\mathscr{E}$, let alone all subsets of $\Omega$ – think of $\Omega$ being the unit interval with Borel $\sigma$-algebra $\mathscr{E}$ and $\mathscr{F}_{t}$ being the set of subsets $E\subseteq[0,1]$ that are countable or have countable complement. So the agent would actually be able to measure the realised scenario without discerning some $E\in\mathscr{E}$, say the interval $[0,\frac{1}{2}]$, although it is a union of measurable singletons. The point is that this union is uncountable, and we do not assume agents to be able to perform uncountable logical operations. Hence, the ability to discern the realised scenario does not necessarily imply the ability to discern all events, that is elements of $\mathscr{E}$. Moreover, all but countably many (and in the case of the Lebesgue measure on $[0,1]$ even all) singletons must have probability zero, and so the ability to discern the realised scenario may not be relevant from a decision-theoretic perspective: If the conditional expectation of some interesting quantity, e.g. a payoff, given $\mathscr{F}_{t}$ is computed, it is completely inconclusive to evaluate it on some $E\in\mathscr{F}_{t}$ whose probability is zero, because the conditional expectation is defined almost surely.
@@ 3.2 Exogenous information structures
Based on the preceding argument about exogenous scenario spaces and filtrations, and bearing in mind the way stochastic decision forests are built on exogenous scenario spaces, the following definition is proposed, which is interpreted and analysed in the sequel.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest on an exogenous scenario space $(\Omega,\mathscr{E})$ and let $\tilde{\bm{\mathsf{X}}}\subseteq{\bm{\mathsf{X}}}$. An exogenous information structure on $\tilde{\bm{\mathsf{X}}}$ is a family $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}}$ such that for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}$, $\mathscr{F}_{\bm{\mathsf{x}}}$ is a sigma-algebra on $D_{\bm{\mathsf{x}}}$ with $\mathscr{F}_{\bm{\mathsf{x}}}\subseteq\mathscr{E}$. An exogenous information structure $\mathscr{F}$ on $\tilde{\bm{\mathsf{X}}}$ is said to admit recall iff for all ${\bm{\mathsf{x}}},{\bm{\mathsf{x}}}^{\prime}\in\tilde{\bm{\mathsf{X}}}$ with ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}^{\prime}$ and every $E\in\mathscr{F}_{\bm{\mathsf{x}}}$, we have $E\cap D_{{\bm{\mathsf{x}}}^{\prime}}\in\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}}$.
@@
Concerning the sense of this definition, we think of any decision maker alias agent as being equipped with some exogenous information structure on the set of his random moves $\tilde{\bm{\mathsf{X}}}$. For any of that agent’s random moves ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}$, $\mathscr{F}_{\bm{\mathsf{x}}}$ is interpreted as the set of events $E\in\mathscr{E}$ relevant at ${\bm{\mathsf{x}}}$, meaning that $E\subseteq D_{\bm{\mathsf{x}}}$, and that the agent can measure at ${\bm{\mathsf{x}}}$, meaning that the agent can discern whether $E$ contains the realised scenario or not. This second property is to be read in the sense discussed in the previous Subsection 3.1. The more straightforward condition of recall [^14] Note that recall of an exogenous information structure is a weak adaptation of the notion of “weak recall” in [42] along the exogenous dimension. ensures that the agent does not forget exogenous information: if the agent is at ${\bm{\mathsf{x}}}^{\prime}$ and has been at ${\bm{\mathsf{x}}}$ already, and was able to measure $E$ at ${\bm{\mathsf{x}}}$, then this agent can measure $E$ given the domain of ${\bm{\mathsf{x}}}^{\prime}$ also at ${\bm{\mathsf{x}}}^{\prime}$. Note that, if for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, we have $D_{\bm{\mathsf{x}}}=\Omega$, then an exogenous information structure on ${\bm{\mathsf{X}}}$ admitting recall is nothing but a monotone decreasing map $\mathscr{F}\colon{\bm{\mathsf{x}}}\mapsto\mathscr{F}_{\bm{\mathsf{x}}}$ from ${\bm{\mathsf{X}}}$ into the set of sub-sigma-algebras of $\mathscr{E}$, “monotone decreasing” meaning that for all ${\bm{\mathsf{x}}},{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ with ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}^{\prime}$, we have $\mathscr{F}_{\bm{\mathsf{x}}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}}$.
@@
It should be noted that what we have at hand here is a generalised adaptation of the notion of filtrations from probability theory to decision trees. In both cases, the index set is partially ordered. However, in the former case the order is total, whereas in the latter quite the opposite case is true, since it is a decision tree. Moreover, the underlying forest can “thin out” towards the crowns non-uniformly across trees alias exogenous scenarios, so that, in general, domains of random moves must be taken into account.